APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday February 2, 2003 (MJD 52673) ---=== OBSERVING TEAM ===--- Mike Harvanek (APO) Atsuko Nitta (APO) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- Most of the night was lost to clouds, wind, dust or some combination thereof. One plate finished (1038) and good data for another plate (1164). ---=== OBSERVING LOG ===--- It was cloudy, windy and dusty at sunset and we remained closed. These conditions, or some combination thereof, remained until approximately 0930Z. However, we were not convinced the conditions were safely stable until 1030Z when we opened to clear skies. Plate 1038 (cartridge 9): FK5 star needed to acquire guide stars. Poor seeing and a telescope attempting to come into thermal equilibrium made focusing difficult. When combined with data from last night (52672), this plate was done in one 30 minute exposure. The 2nd exposure was aborted when we discovered we had finished the plate with the first exposure. Plate 1164 (cartridge 8): FK5 star needed to acquire guide stars. One good 20 minute exposure. The second exposure was cut short at 12 minutes due to twilight and the seeing was sufficiently poor that this exposure was not useful for 3 of the 4 cameras. There were more than a thousand MCP watcher errors like the following: MJD CONT id = 3 djm = 52673 (2003-02-03 09:04:22Z sdssmcp) (MJD CONT id = 3 djm = 52673 (2003-02-03 09:04:22Z sdssmcp): 2003-02-03 09:04:35Z ) and one of these: rotator ((throttle) MJD CONT id = 3 djm = 52673 (2003-02-03 12:09:29Z sdssmcp): 2003-02-03 12:09:41Z) It appears to be getting harder to get the southwest pin in the building. After hitting the stop with what I thought was plenty sufficient force, I still could not get the southwest pin to go into the second (lower) hole. I had to back the building up and try it again. This time I succeeded in getting the pin in but I overloaded the building motor box in the process. I cleared the overload. Can we bore out the southwest pin holes a little? ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 23:31 19069 0.0 bias ----- sequence 19070, plate 1038 ------- 23:35 19070 2.1 arc for collimation check 23:38 19071 2.1 arc " " " 00:14 19072 2.1 arc " " " 00:17 19073 2.1 arc " " " 06:52 19074 0.0 dark for testing PR 06:54 19075 0.0 dark " " " 06:57 19076 5.0 dark " " " 07:04 19077 5.0 dark " " " 07:06 19078 5.0 dark " " " 11:12 19079 10.0 flat 11:14 19080 2.0 arc 11.7 8.4 9.8 7.5 tonight 11:51 19081 1800.1 target 6.9 6.1 5.6 5.7 52672 12:00 19082 87.0 target aborted ------------------- 12:08 19083 240.0 smear 18.6 14.5 15.4 13.2 DONE ----- sequence 19084, plate 1164 ------- 12:44 19084 1200.1 target 12:59 19085 712.3 target not useful 13:02 19086 10.0 flat 13:04 19087 2.0 arc 7.9 5.7 6.5 5.0 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 11:29Z 9 1038 192.00 -0.0067 67.26 0.0011 17.62 -0.0027 1.000130 12:24Z 8 1164 161.27 -0.0046 68.12 -0.0009 -26.72 0.0000 1.000310 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL6000 Stays: JL6001 ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 11:30Z 9 1038 1.00013 -1283 424 350 192 67.3 -4.3 11 325 ---- 2.5" 12:23Z 8 1164 1.00026 -2566 -690 300 161 68.1 -4.7 15 340 ---- 2.4" ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:18Z 44 24 41 257 (WSW) 4701 - cloudy, windy, dusty 23:49Z 43 23 30 263 (W) 6494 - " " " 00:20Z 43 22 27 244 (WSW) 6557 - " " " 00:51Z 42 21 28 258 (WSW) 6989 - " " " 01:21Z 42 17 28 266 (W) 13611 - " " " 01:52Z 41 16 30 256 (WSW) 12453 - " " " 02:22Z 39 13 24 263 (W) 10736 - " " " 02:52Z 39 12 27 273 (W) 5275 - " " " 03:25Z 38 10 29 268 (W) 3647 - pcl, " " 03:58Z 37 11 17 243 (WSW) 3080 - pcl, " " 04:29Z 36 12 33 259 (W) 2260 - clear, " 05:02Z 35 11 25 268 (W) 2616 - clear, " 05:33Z 34 11 16 273 (W) 5409 - clear, " " 06:06Z 32 10 28 285 (WNW) 2732 - clear, " " 06:38Z 31 7 19 289 (WNW) 2661 - pcl, " " 07:10Z 30 7 23 284 (WNW) 3401 - pcl, " " 07:42Z 30 5 25 268 (W) 6532 - pcl, " 08:14Z 28 4 24 253 (WSW) 7028 - pcl, " 08:47Z 28 4 20 248 (WSW) 2303 - pcl, " " 09:19Z 27 0 13 267 (W) 2343 - pcl 09:51Z 27 0 17 270 (W) 1940 - clear 10:23Z 26 0 20 273 (W) 1238 - clear 10:56Z 25 0 11 267 (W) 1587 - clear 11:27Z 24 1 14 318 (NW) 1362 - clear 11:57Z 24 0 12 311 (NW) 1465 - clear 12:28Z 23 1 18 331 (NNW) 1731 - clear 13:02Z 22 1 16 329 (NNW) 1747 - clear pcl = partly cloudy 13:33Z 22 2 15 318 (NW) 1677 - ---=== TELESCOPE STATUS ===--- 10:30Z Enclosure doors opened, fans and blowers on, louvers opened. 10:35Z Enclosure off. 13:05Z Enclosure on, doors closed, fans and blowers off, louvers closed. Telescope stowed at 30 degrees with cartridge 8 (plate 1164) mounted. Counterweights at 280. Spectro auto-fill connected. Imager dewar 13:25 65 lb Spectro dewar 13:25 109 lb No interlocks bypassed. ---=== SOFTWARE USED ===--- IOP/SOP: v3_105_0 Watcher: v2_20_0 MCP: v5_17_0 TPM: tpm_v2_24_0 AstroDa: v14_47 TCC: TCC 2.6.8 sdssProcedures: v1_62 tccMon: v1_9 SoS: v4_9_13 hoggPT: v1_5_8 plate-mapper: v4_2_0 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.0820 0.0790 Axial B 0.7990 0.7970 Axial C 0.8110 0.8240 Trans D -9.0790 -9.1130 Lateral E 1.8510 1.8870 Lateral F 0.0000 1.4300 GALILS Commanded: 5400. -3700. 900. -200. 31550. 30650. Actual: 5400. -3700. 900. -200. 31550. 30650. SETMIR VALUES PriDesOrient: 0.00 -11.80 23.00 1256.90 642.10 PriOrient: 0.00 -12.16 22.81 1257.41 642.19 SECONDARY: ---------- Focus: 0.00 Air Temp.: 6.1 Alt.: 29.997904 MIGS TONIGHT NOMINAL Axial A 1.5170 1.5170 Axial B 1.1560 1.1560 Axial C 1.1000 1.0990 Trans D -0.6890 -0.7070 GALILS Commanded: 1619706. 1555189. 1576336. -3400. -7900. Actual: 1619650. 1555250. 1576350. -3400. -7900. SETMIR VALUES SecDesOrient: 1257.00 0.00 -20.00 0.00 130.98 SecOrient: 1257.00 -0.01 -20.04 -0.51 130.89 ---=== PROBLEMS IN DETAIL ===---